# Decentralized Stochastic Subgradient Methods for Nonsmooth Nonconvex Optimization

### v1

@@ 3.1 Basic assumptions
    ${\bm{W}}\in\mathbb{R}^{d\times d}$ is a mixing matrix.
@@
    The sequence $\{{\bm{Z}}_{k}\}$ and $\{{\bm{H}}_{k}\}$ are bounded.
@@
    The sequences of step-sizes $\{\eta_{k}\}$ is positive and diminishing, which satisfies $\sum_{k=0}^{+\infty}\eta_{k}=+\infty$. Moreover, for any $T>0$, it holds that
    $\lim_{i\to+\infty}\sup_{i\leq j\leq\Lambda(\lambda(i)+T)}\left\|\sum_{k=i}^{j}\eta_{k}\Xi_{k+1}\right\|=0.$ (3.1)
  4. 4.
    There exists a positive diminishing sequence $\{\delta_{k}\}$ and a locally bounded graph-closed set-valued mapping $\mathcal{H}_{i}$, such that for each $i\in[d]$, it holds that
    ${\bm{H}}_{k}{\bm{e}}\in\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}^{\delta_{k}}({\bm{z}}_{i,k})\right),\quad\forall k\in\mathbb{N}.$ (3.2)
Here we present some comments on Assumption 3.1. Assumption 3.1(2) assumes the uniform boundedness of the sequence $\{{\bm{Z}}_{k}\}$, which is commonly employed in existing works [8, 15, 46, 45]. Assumption 3.1(3) illustrates that the noises in the framework (DSM) can be controlled by the step-sizes, which is identical to the condition on the noises in [2]. Furthermore, Assumption 3.1(4) illustrates that the averaged update directions of all the agents $\{{\bm{H}}_{k}{\bm{e}}\}$ approximate the evaluation of $\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}({\bm{z}}_{i,k})$. For conciseness, we define the set-valued mapping $\mathcal{H}:\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ as
$\mathcal{H}({\bm{z}}):=\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}({\bm{z}})\right).$ (3.3)
Then we present the following assumptions on the differential inclusion
$\frac{\mathrm{d}{\bm{z}}}{\mathrm{d}t}\in-\mathcal{H}({\bm{z}}),$ (3.4)
which corresponds to the discrete update scheme in framework (DSM).
@@
Assumption 3.2(1) clarifies the Lyapunov properties of a differential inclusion with respect to $-\mathcal{H}$, which is standard in the literature and can be satisfied by most appropriately selected set-valued mappings. Assumption 3.2(2) is the Weak Sard’s theorem, which stipulates that the set of values of $\Psi$ at points outside the stable set is dense in $\mathbb{R}$. As demonstrated in demonstrated in [15, 11], Assumption 3.2(2) is mild in practice.
@@ 3.2 Global convergence
In this subsection, we establish the convergence properties of our proposed framework (DSM), under Assumption 3.1 and Assumption 3.2. Towards this end, we first conduct consensus analysis of $\{{\bm{Z}}_{k}\}$. Denote
${\bm{P}}={\bm{e}}{\bm{1}}_{d}^{\top},\quad{\bm{P}}_{\perp}={\bm{I}}_{d}-{\bm{e}}{\bm{1}}_{d}^{\top},$
then any matrix ${\bm{Z}}\in\mathbb{R}^{m\times d}$ has the following orthogonal decomposition,
${\bm{Z}}={\bm{Z}}{\bm{P}}+{\bm{Z}}{\bm{P}}_{\perp}.$ (3.5)
Define the averaging vector $\hat{{\bm{Z}}}:={\bm{Z}}{\bm{e}}$, and the disagreement matrix ${\bm{Z}}_{\perp}:={\bm{Z}}{\bm{P}}_{\perp}$. Then (3.5) can be rephrased as
${\bm{Z}}=\hat{{\bm{Z}}}{\bm{1}}_{d}^{\top}+{\bm{Z}}_{\perp}$ (3.6)
We begin our proof with the following preliminary lemma from [40, Lemma 3.1].
@@
Let $\{\gamma_{k}\}$ be a sequence of real numbers. If $10<\alpha<$ 1 and $\lim_{k\rightarrow+\infty}\gamma_{k}=\gamma$, then it holds that
$\lim_{k\rightarrow+\infty}\sum_{\ell=0}^{k}\alpha^{k-\ell}\gamma_{\ell}=\frac{\gamma}{1-\alpha}.$
Next, we present the following lemma to show that $\left\|({\bm{Z}}_{k+1}-{\bm{Z}}_{k}){\bm{e}}\right\|$ converges to 00 as $k$ goes to infinity.
@@
Suppose Assumption 3.1 holds, then for any $\{{\bm{Z}}_{k}\}$ generated by the framework (DSM), we have $\lim_{k\to+\infty}\left\|\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})\right\|=0$. Moreover, it holds that
$\lim_{k\to+\infty}\left\|({\bm{Z}}_{k+1}-{\bm{Z}}_{k}){\bm{e}}\right\|=0.$ (3.7)
@@
$\limsup_{i\to+\infty}\left\|\eta_{i}\Xi_{i+1}\right\|\leq\lim_{i\to+\infty}\sup_{i\leq j\leq\Lambda(\lambda(i)+T)}\left\|\sum_{k=i}^{j}\eta_{k}\Xi_{k+1}\right\|=0.$ (3.8)
@@
Moreover, from the update scheme of $\{{\bm{Z}}_{k}\}$ in (DSM) and Definition 2.1, it holds that
${\bm{Z}}_{k+1}{\bm{e}}={\bm{Z}}_{k}{\bm{e}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1}){\bm{e}}.$ (3.10)
Together with (3.9), it holds that
$\lim_{k\to+\infty}\left\|({\bm{Z}}_{k+1}-{\bm{Z}}_{k}){\bm{e}}\right\|=\lim_{k\to+\infty}\left\|\eta_{k}({\bm{H}}_{k}+\Xi_{k+1}){\bm{e}}\right\|\leq\lim_{k\to+\infty}\frac{1}{\sqrt{d}}\left\|\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})\right\|=0.$ (3.11)
@@
Next, we denote ${\bm{Z}}_{\perp,k}:={\bm{Z}}_{k}{\bm{P}}_{\perp}$ as the disagreement matrix at $k$-th iteration. Intuitively, ${\bm{Z}}_{\perp,k}$ measures the dissimilarity of the local variables of all the agents at $k$-th iteration. In the following proposition, we prove that $\lim_{k\to+\infty}{\bm{Z}}_{\perp,k}=0$, i.e., the dissimilarity of local variables of all the agents converges to 00.
@@
$\left\|{\bm{Z}}_{\perp,k+1}\right\|=\left\|[{\bm{Z}}_{k}{\bm{W}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})]({\bm{I}}_{d}-{\bm{P}})\right\|$ (3.13)
@@
Here the second inequality follows from the fact that $({\bm{I}}_{d}-{\bm{P}}){\bm{W}}({\bm{I}}_{d}-{\bm{P}})={\bm{W}}({\bm{I}}_{d}-{\bm{P}})$. Since Lemma 3.4 illustrates that $\lim_{k\to+\infty}\left\|\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})\right\|=0$, then it can be readily deduced by Lemma 3.3 that
$\lim_{k\to+\infty}\left\|{\bm{Z}}_{\perp,k}\right\|=0.$ (3.14)
This completes the proof. ∎
As illustrated in Proposition 3.5, any sequence $\{{\bm{Z}}_{k}\}$ from (DSM) satisfies
$\lim_{k\to+\infty}\left\|{\bm{Z}}_{k}-{\bm{Z}}_{k}{\bm{P}}\right\|=\lim_{k\to+\infty}\left\|{\bm{Z}}_{k}-{\bm{Z}}_{k}{\bm{e}}{\bm{1}}_{d}^{\top}\right\|=0.$ (3.15)
As a result, we can conclude that the cluster points of the sequence $\{{\bm{Z}}_{k}\}$ are exactly the same as those of the sequence $\{{\bm{Z}}_{k}{\bm{e}}{\bm{1}}_{d}^{\top}\}$. To analyze the convergence properties of $\{{\bm{Z}}_{k}{\bm{e}}\}$, we present the following proposition to illustrate the update scheme of $\{{\bm{Z}}_{k}{\bm{e}}\}$.
@@
${\bm{Z}}_{k+1}{\bm{e}}\in{\bm{Z}}_{k}{\bm{e}}-\eta_{k}\left(\mathcal{H}^{\tilde{\delta}_{k}}({\bm{Z}}_{k}{\bm{e}})+\Xi_{k+1}{\bm{e}}\right).$ (3.16)
@@
${\bm{Z}}_{k+1}{\bm{e}}={\bm{Z}}_{k}{\bm{W}}{\bm{e}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1}){\bm{e}}={\bm{Z}}_{k}{\bm{e}}-\eta_{k}{\bm{H}}_{k}{\bm{e}}-\eta_{k}\Xi_{k+1}{\bm{e}}.$ (3.17)
Let $\delta^{\star}_{k}=\left\|{\bm{Z}}_{k}({\bm{I}}_{d}-{\bm{P}})\right\|$, from the definition of ${\bm{P}}$, it holds that $\left\|{\bm{z}}_{i,k}-{\bm{Z}}_{k}{\bm{e}}\right\|\leq\left\|{\bm{Z}}_{k}({\bm{I}}_{d}-{\bm{P}})\right\|=\delta^{\star}_{k}$. Together with Assumption 3.1(3), we can conclude that
${\bm{H}}_{k}{\bm{e}}\in\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}^{\delta_{k}}({\bm{z}}_{i,k})\right)\subseteq\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}^{\delta_{k}+\delta^{\star}_{k}}({\bm{Z}}_{k}{\bm{e}})\right).$ (3.18)
@@
$\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}^{\delta_{k}+\delta^{\star}_{k}}({\bm{Z}}_{k}{\bm{e}})\right)\subseteq\mathcal{H}^{\tilde{\delta}_{k}}({\bm{Z}}_{k}{\bm{e}}).$ (3.19)
@@
$\sup\left\{\mathrm{dist}\left({\bm{y}},\mathcal{H}^{\delta_{\varepsilon}}({\bm{Z}}_{k_{j}}{\bm{e}})\right):{\bm{y}}\in\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}^{\delta_{k_{j}}+\delta^{\star}_{k_{j}}}({\bm{Z}}_{k_{j}}{\bm{e}})\right)\right\}>0.$ (3.20)
From the uniform boundedness of $\{{\bm{Z}}_{k}\}$ in Assumption 3.1(1), without loss of generality, we can assume that $\{{\bm{Z}}_{k_{j}}\}$ converges to $\tilde{{\bm{Z}}}\in\mathbb{R}^{n\times d}$. Then from the graph closedness and local boundedness of $\mathcal{H}_{i}$, we have
$\lim_{j\to+\infty}\sup\left\{\mathrm{dist}\left({\bm{y}},\mathcal{H}_{i}(\tilde{{\bm{Z}}}{\bm{e}})\right):{\bm{y}}\in\mathcal{H}_{i}^{\delta_{k_{j}}+\delta^{\star}_{k_{j}}}({\bm{Z}}_{k_{j}}{\bm{e}})\right\}=0.$ (3.21)
@@
$\lim_{j\to+\infty}\sup\left\{\mathrm{dist}\left({\bm{y}},\mathcal{H}(\tilde{{\bm{Z}}}{\bm{e}})\right):{\bm{y}}\in\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\mathcal{H}_{i}^{\delta_{k_{j}}+\delta^{\star}_{k_{j}}}({\bm{Z}}_{k_{j}}{\bm{e}})\right)\right\}=0.$ (3.22)
@@
${\bm{Z}}_{k+1}{\bm{e}}={\bm{Z}}_{k}{\bm{W}}{\bm{e}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1}){\bm{e}}\in{\bm{Z}}_{k}{\bm{e}}-\eta_{k}\left(\mathcal{H}^{\tilde{\delta}_{k}}({\bm{Z}}_{k}{\bm{e}})+\Xi_{k+1}{\bm{e}}\right).$ (3.23)
@@
Suppose Assumption 3.1 holds. Then the interpolated process of the sequence $\{{\bm{Z}}_{k}{\bm{e}}\}$ is a perturbed solution to the differential inclusion (3.4).
@@
Firstly, from Proposition 3.6 we can conclude that there exists a nonnegative diminishing sequence $\{\tilde{\delta}_{k}\}$ such that the sequence $\{{\bm{Z}}_{k}{\bm{e}}\}$ follows the update scheme
${\bm{Z}}_{k+1}{\bm{e}}={\bm{Z}}_{k}{\bm{W}}{\bm{e}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1}){\bm{e}}\in{\bm{Z}}_{k}{\bm{e}}-\eta_{k}\left(\mathcal{H}^{\tilde{\delta}_{k}}({\bm{Z}}_{k}{\bm{e}})+\Xi_{k+1}{\bm{e}}\right).$ (3.24)
@@
$\lim_{i\to+\infty}\sup_{i\leq j\leq\Lambda(\lambda(i)+T)}\left\|\sum_{k=i}^{j}\eta_{k}\Xi_{k+1}{\bm{e}}\right\|\leq\frac{1}{\sqrt{d}}\lim_{i\to+\infty}\sup_{i\leq j\leq\Lambda(\lambda(i)+T)}\left\|\sum_{k=i}^{j}\eta_{k}\Xi_{k+1}\right\|=0.$ (3.25)
Then combined with Assumption 3.1(2), it directly follows from Lemma 2.11 that the interpolated process of the sequence $\{{\bm{Z}}_{k}{\bm{e}}\}$ is a perturbed solution of the differential inclusion (3.4). This completes the proof. ∎
Based on Proposition 3.7 and Theorem 2.12, we establish the convergence theorem, where we prove that the sequence $\{{\bm{Z}}_{k}\}$ converges to the stable set of (3.4) for our framework (DSM).
@@
Proposition 3.7 illustrates that the interpolated process of $\{{\bm{Z}}_{k}{\bm{e}}\}$ is a perturbed solution to the differential inclusion (3.4). Together with Assumption 3.2 and Theorem 2.12, we can conclude that any cluster point of $\{{\bm{Z}}_{k}{\bm{e}}\}$ lies in $\mathcal{A}$, and the sequence of function values $\{\Psi({\bm{Z}}_{k}{\bm{e}})\}$ converges. Moreover, notice that Proposition 3.5 illustrates that
$\lim_{k\to+\infty}\left\|{\bm{Z}}_{k}-{\bm{Z}}_{k}{\bm{e}}{\bm{1}}_{d}^{\top}\right\|=0.$ (3.26)
Therefore, we can conclude that any cluster point of $\{{\bm{Z}}_{k}\}$ lies in $\{{\bm{Z}}\in\mathbb{R}^{m\times d}:{\bm{Z}}={\bm{Z}}{\bm{1}}_{d}^{\top},~{}{\bm{Z}}\in\mathcal{A}\}$, and the sequence of function values $\{\Psi({\bm{z}}_{i,k})\}$ converges for each $i\in[d]$. ∎
4 Application: Decentralized Stochastic Subgradient Methods
In this section, we aim to illustrate the great flexibility of our proposed framework (DSM). We show that our proposed framework (DSM) encloses a wide range of popular decentralized stochastic subgradient methods, including DSGD, DSGD with momentum (DSGDm), and DSGD with gradient-tracking technique (DSGD-T). More importantly, we demonstrate that our theoretical results in Theorem 3.8 provide theoretical convergence guarantees for these decentralized stochastic subgradient methods, especially in the minimization of nonsmooth path-differentiable functions. Throughout this section, we make the following assumptions on (DOP).

### v2

@@ 3.1 Basic assumptions
    The sequences $\{{\bm{Z}}_{k}\}$ and $\{{\bm{H}}_{k}\}$ are bounded.
@@
    The sequence of step-sizes $\{\eta_{k}\}$ is positive and diminishing, which satisfies $\sum_{k=0}^{+\infty}\eta_{k}=+\infty$. Moreover, for any $T>0$, it holds that
    $\lim_{i\to+\infty}\sup_{i\leq j\leq\Lambda_{\eta}(\lambda_{\eta}(i)+T)}\left\|\sum_{k=i}^{j}\eta_{k}\Xi_{k+1}\right\|=0.$
@@
    There exists a positive diminishing sequence $\{\delta_{k}\}$ and a locally bounded graph-closed set-valued mapping $\Phi_{i}$, such that for each $i\in[d]$, it holds that
    $\frac{1}{d}{\bm{H}}_{k}{\bm{1}_{d}}\in\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}^{\delta_{k}}({\bm{z}}_{i,k})\right),\quad\forall k\in\mathbb{N}.$ (3.1)
Remark 3.2.
Assumption 3.1(1) assumes the uniform boundedness of the sequence $\{{\bm{Z}}_{k}\}$, which is commonly employed in existing works [28, 29, 36, 39]. Assumption 3.1(2) illustrates that the noises in the framework (DSM) can be controlled by the step-sizes, which is identical to the condition on the noises in [30]. Furthermore, Assumption 3.1(3) illustrates that the averaged update directions of all the agents $\{\frac{1}{d}{\bm{H}}_{k}{\bm{1}_{d}}\}$ approximate the evaluation of $\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}({\bm{z}}_{i,k})$.
For conciseness, we define the set-valued mapping $\Phi:\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ as $\Phi({\bm{z}}):=\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}({\bm{z}})\right)$. Then we present the following assumptions on the differential inclusion
$\frac{\mathrm{d}{\bm{z}}}{\mathrm{d}t}\in-\Phi({\bm{z}}),$ (3.2)
which corresponds to the discrete update scheme in (DSM).
@@
Assumption 3.3(1) clarifies the Lyapunov properties of a differential inclusion with respect to $-\Phi$, which is standard in the literature and can be satisfied by most appropriately selected set-valued mappings. Assumption 3.3(2) is the Weak Sard’s theorem, which stipulates that the set of values of $\psi$ at points outside the stable set is dense in $\mathbb{R}$. As demonstrated in demonstrated in [29, 37], Assumption 3.3(2) is mild in practice.
@@ 3.2 Global convergence
In this subsection, we establish the convergence properties of our proposed framework (DSM), under Assumption 3.1 and Assumption 3.3.
Towards this end, we first conduct consensus analysis of $\{{\bm{Z}}_{k}\}$. Denote ${\bm{P}}=\frac{1}{d}{\bm{1}_{d}}{\bm{1}}_{d}^{\top}$ and ${\bm{P}}_{\perp}={\bm{I}}_{d}-\frac{1}{d}{\bm{1}_{d}}{\bm{1}}_{d}^{\top}$, then any matrix ${\bm{Z}}_{k}\in\mathbb{R}^{m\times d}$ has the orthogonal decomposition ${\bm{Z}}_{k}={\bm{Z}}_{k}{\bm{P}}+{\bm{Z}}_{k}{\bm{P}}_{\perp}$. Define the disagreement matrix ${\bm{Z}}_{\perp,k}:={\bm{Z}}_{k}{\bm{P}}_{\perp}$. Then the above orthogonal decomposition can be rephrased as ${\bm{Z}}_{k}={\bm{Z}}_{k}{\bm{P}}+{\bm{Z}}_{\perp,k}$.
We begin our proof with the following preliminary lemma from [53, Lemma 3.1].
@@
Let $\{\gamma_{k}\}$ be a sequence of real numbers. If $10<\alpha<$ 1 and $\lim_{k\rightarrow+\infty}\gamma_{k}=\gamma$, then it holds that $\lim_{k\rightarrow+\infty}\sum_{\ell=0}^{k}\alpha^{k-\ell}\gamma_{\ell}=\frac{\gamma}{1-\alpha}$.
Next, we present the following lemma to show that $\left\|\frac{1}{d}({\bm{Z}}_{k+1}-{\bm{Z}}_{k}){\bm{1}}_{d}\right\|$ converges to 00 as $k$ goes to infinity.
@@
Suppose Assumption 3.1 holds, then for any $\{{\bm{Z}}_{k}\}$ generated by the framework (DSM), we have $\lim_{k\to+\infty}\left\|\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})\right\|=0$. Moreover, it holds that $\lim_{k\to+\infty}\left\|\frac{1}{d}({\bm{Z}}_{k+1}-{\bm{Z}}_{k}){\bm{1}_{d}}\right\|=0$.
@@
$\limsup_{i\to+\infty}\left\|\eta_{i}\Xi_{i+1}\right\|\leq\lim_{i\to+\infty}\sup_{i\leq j\leq\Lambda_{\eta}(\lambda_{\eta}(i)+T)}\left\|\sum_{k=i}^{j}\eta_{k}\Xi_{k+1}\right\|=0.$
@@
Moreover, from the update scheme of $\{{\bm{Z}}_{k}\}$ in (DSM) and Definition 2.1, it holds that $\frac{1}{d}{\bm{Z}}_{k+1}{\bm{1}_{d}}=\frac{1}{d}{\bm{Z}}_{k}{\bm{1}_{d}}-\frac{\eta_{k}}{d}({\bm{H}}_{k}+\Xi_{k+1}){\bm{1}_{d}}$. Together with (3.3), we have
$\lim_{k\to+\infty}\frac{1}{d}\left\|({\bm{Z}}_{k+1}-{\bm{Z}}_{k}){\bm{1}}_{d}\right\|=\lim_{k\to+\infty}\left\|\frac{\eta_{k}}{d}({\bm{H}}_{k}+\Xi_{k+1}){\bm{1}}_{d}\right\|$
[…]
@@
Intuitively, ${\bm{Z}}_{\perp,k}$ measures the dissimilarity of local variables of all the agents at $k$-th iteration. In the following proposition, we prove that $\lim_{k\to+\infty}{\bm{Z}}_{\perp,k}=0$, i.e., the dissimilarity of local variables of all the agents converges to 00.
@@
$\left\|{\bm{Z}}_{\perp,k+1}\right\|=\left\|[{\bm{Z}}_{k}{\bm{W}}-\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})]({\bm{I}}_{d}-{\bm{P}})\right\|$
@@
Here the second inequality follows from the fact that $({\bm{I}}_{d}-{\bm{P}}){\bm{W}}({\bm{I}}_{d}-{\bm{P}})={\bm{W}}({\bm{I}}_{d}-{\bm{P}})$. Since Lemma 3.5 illustrates that $\lim_{k\to+\infty}\left\|\eta_{k}({\bm{H}}_{k}+\Xi_{k+1})\right\|=0$, then it can be readily deduced by Lemma 3.4 that $\lim_{k\to+\infty}\left\|{\bm{Z}}_{\perp,k}\right\|=0$. This completes the proof. ∎
As illustrated in Proposition 3.6, any sequence $\{{\bm{Z}}_{k}\}$ from (DSM) satisfies $\lim_{k\to+\infty}\left\|{\bm{Z}}_{k}-{\bm{Z}}_{k}{\bm{P}}\right\|=0$. As a result, we can conclude that the cluster points of the sequence $\{{\bm{Z}}_{k}\}$ are exactly the same as those of the sequence $\{{\bm{Z}}_{k}{\bm{P}}\}$.
To analyze the convergence properties of $\{\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\}$, we present the following proposition to illustrate the update scheme of $\{\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\}$.
@@
$\frac{1}{d}{\bm{Z}}_{k+1}{\bm{1}}_{d}\in\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}-\eta_{k}\left(\Phi^{\tilde{\delta}_{k}}(\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d})+\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d}\right).$ (3.5)
@@
$\frac{1}{d}{\bm{Z}}_{k+1}{\bm{1}}_{d}$ $=\frac{1}{d}{\bm{Z}}_{k}{\bm{W}}{\bm{1}}_{d}-\frac{\eta_{k}}{d}({\bm{H}}_{k}+\Xi_{k+1}){\bm{1}}_{d}$
$=\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}-\frac{\eta_{k}}{d}{\bm{H}}_{k}{\bm{1}}_{d}-\frac{\eta_{k}}{d}\Xi_{k+1}{\bm{1}}_{d}.$
Let $\delta^{\star}_{k}=\left\|{\bm{Z}}_{k}{\bm{P}}_{\perp}\right\|$, from the definition of ${\bm{P}}_{\perp}$, it holds that $\left\|{\bm{z}}_{i,k}-\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\right\|\leq\left\|{\bm{Z}}_{k}{\bm{P}}_{\perp}\right\|=\delta^{\star}_{k}$.
For simplicity, we denote $\mathrm{conv}\,(\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}^{\delta_{k}+\delta^{\star}_{k}}(\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}))$ by $\mathcal{C}_{k}$. Together with Assumption 3.1(3), we can conclude that
$\frac{1}{d}{\bm{H}}_{k}{\bm{1}}_{d}\in\mathrm{conv}\,\left(\frac{1}{d}\sum_{i=1}^{d}\Phi_{i}^{\delta_{k}}({\bm{z}}_{i,k})\right)\subseteq\mathcal{C}_{k}.$ (3.6)
@@
$\mathcal{C}_{k}\subseteq\Phi^{\tilde{\delta}_{k}}(\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}).$ (3.7)
@@
$\sup\left\{\mathrm{dist}\left({\bm{y}},\Phi^{\delta_{\varepsilon}}(\frac{1}{d}{\bm{Z}}_{k_{j}}{\bm{1}}_{d})\right):{\bm{y}}\in\mathcal{C}_{k_{j}}\right\}>0.$ (3.8)
From the uniform boundedness of $\{{\bm{Z}}_{k}\}$ in Assumption 3.1(1), without loss of generality, we can assume that $\{{\bm{Z}}_{k_{j}}\}$ converges to $\tilde{{\bm{Z}}}\in\mathbb{R}^{n\times d}$. Then from the graph closedness and local boundedness of $\Phi_{i}$, we have
$\lim_{j\to+\infty}\sup\left\{\mathrm{dist}\left({\bm{y}},\Phi_{i}(\frac{\tilde{{\bm{Z}}}{\bm{1}}_{d}}{d})\right):{\bm{y}}\in\Phi_{i}^{\delta_{k_{j}}+\delta^{\star}_{k_{j}}}(\frac{{\bm{Z}}_{k_{j}}{\bm{1}}_{d}}{d})\right\}=0.$
@@
$\lim_{j\to+\infty}\sup\left\{\mathrm{dist}\left({\bm{y}},\Phi(\frac{1}{d}\tilde{{\bm{Z}}}{\bm{1}}_{d})\right):{\bm{y}}\in\mathcal{C}_{k_{j}}\right\}=0.$
@@
$\frac{1}{d}{\bm{Z}}_{k+1}{\bm{1}}_{d}$ $=\frac{1}{d}{\bm{Z}}_{k}{\bm{W}}{\bm{1}}_{d}-\eta_{k}({\bm{H}}_{k}+\frac{1}{d}\Xi_{k+1}){\bm{1}}_{d}$ (3.9)
$\in\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}-\eta_{k}\left(\Phi^{\tilde{\delta}_{k}}(\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d})+\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d}\right).$
@@
Suppose Assumption 3.1 holds. Then the interpolated process of the sequence $\{\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\}$ is a perturbed solution to the differential inclusion (3.2).
@@
Firstly, from Proposition 3.7 we can conclude that there exists a nonnegative diminishing sequence $\{\tilde{\delta}_{k}\}$ such that the sequence $\{\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\}$ follows the update scheme
$\frac{1}{d}{\bm{Z}}_{k+1}{\bm{1}}_{d}$ $=\frac{1}{d}{\bm{Z}}_{k}{\bm{W}}{\bm{1}}_{d}-\eta_{k}\frac{1}{d}({\bm{H}}_{k}+\Xi_{k+1}){\bm{1}}_{d}$ (3.10)
$\in\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}-\eta_{k}\left(\Phi^{\tilde{\delta}_{k}}(\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d})+\frac{1}{d}\Xi_{k+1}{\bm{1}}_{d}\right).$
@@
$\lim_{i\to+\infty}\sup_{i\leq j\leq\Lambda_{\eta}(\lambda_{\eta}(i)+T)}\left\|\sum_{k=i}^{j}\frac{\eta_{k}}{d}\Xi_{k+1}{\bm{1}}_{d}\right\|$ (3.11)
[…]
Then combined with Assumption 3.1(1), it directly follows from [36, Lemma 2.20] that the interpolated process of the sequence $\{\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\}$ is a perturbed solution of the differential inclusion (3.2). This completes the proof.
Based on Proposition 3.8 and Theorem 2.10, we establish the convergence theorem, where we prove that $\{{\bm{z}}_{i,k}\}$ converges subsequentially to the stable set of (3.2) for our framework (DSM).
@@
Proposition 3.8 illustrates that the interpolated process of $\{\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\}$ is a perturbed solution to the differential inclusion (3.2). Together with Assumption 3.3 and Theorem 2.10, we can conclude that any cluster point of $\{\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d}\}$ lies in $\mathcal{A}$, and the sequence of function values $\{\psi(\frac{1}{d}{\bm{Z}}_{k}{\bm{1}}_{d})\}$ converges.
Moreover, notice that Proposition 3.6 illustrates that $\lim_{k\to+\infty}\left\|{\bm{Z}}_{k}-{\bm{Z}}_{k}{\bm{P}}\right\|=0$. Therefore, we can conclude that any cluster point of $\{{\bm{Z}}_{k}\}$ lies in $\{{\bm{Z}}\in\mathbb{R}^{m\times d}:{\bm{Z}}={\bm{z}}{\bm{1}}_{d}^{\top},~{}{\bm{z}}\in\mathcal{A}\}$, and the sequence of function values $\{\psi({\bm{z}}_{i,k})\}$ converges for each $i\in[d]$.
4 Application: Decentralized Stochastic Subgradient-based Methods
In this section, we aim to illustrate the great flexibility of our proposed framework (DSM). We show that our proposed framework (DSM) encloses a wide range of popular decentralized stochastic subgradient-based methods, including DSGD, DSGD with momentum (DSGD-M), and DSGD with gradient-tracking technique (DSGD-T). More importantly, we demonstrate that our theoretical results in Theorem 3.9 provide theoretical convergence guarantees for these decentralized stochastic subgradient-based methods, especially in the minimization of nonsmooth path-differentiable functions.
Throughout this section, we make the following assumptions on (DOP).